$SPAD/src/input unittest3.input

Timothy Daly December 3, 2016

Abstract

Unit test the user level commands

Contents

2

— * —

)set break resume
)sys rm -f unittest3.output
)spool unittest3.output
)set mes auto off
)clear all
--S 1 of 75
)lisp (identity |$inputPromptType|)
--R
--RValue = |step|
--E 1
--S 2 of 75
)lisp (setq |$inputPromptType| '|none|)
--R
--RValue = |none|
--E 2
--S 3 of 75
1
--R   (1)  1
--R                                                        Type: PositiveInteger
--E 3
--S 4 of 75
)lisp (setq |$inputPromptType| '|plain|)
--RValue = |plain|
--E 4
--S 5 of 75
2
--R
--R   (2)  2
--R                                                        Type: PositiveInteger
--E 5
--S 6 of 75
)lisp (setq |$inputPromptType| '|step|)
--R
--RValue = |step|
--E 6
--S 7 of 75
2
--R
--R   (3)  2
--R                                                        Type: PositiveInteger

--E 7
--S 8 of 75
)lisp (setq |$inputPromptType| '|frame|)
--R
--RValue = |frame|
--E 8
--S 9 of 75
2
--R
--R   (4)  2
--R                                                        Type: PositiveInteger
--E 9
--S 10 of 75
)lisp (setq |$inputPromptType| t)
--R
--RValue = T
--E 10
--S 11 of 75
2
--R
--R   (5)  2
--R                                                        Type: PositiveInteger
--E 11
--S 12 of 75
)set debug
--R                    Current Values of  debug  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rlambdatype   show type information for #1 syntax        off
--Rdalymode     Interpret leading open paren as lisp       off
--R
--E 12
--S 13 of 75
)set debug lambdatype
--R-------------------------- The lambdatype Option --------------------------
--R
--R Description: show type information for #1 syntax
--R
--R The lambdatype option may be followed by any one of the following:
--R
--R    on
--R -> off
--R

--R The current setting is indicated.
--R
--E 13
--S 14 of 75
)set debug lambdatype on
--E 14
--S 15 of 75
)set debug lambdatype
--R-------------------------- The lambdatype Option --------------------------
--R
--R Description: show type information for #1 syntax
--R
--R The lambdatype option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 15
--S 16 of 75
)set debug dalymode
--R--------------------------- The dalymode Option ---------------------------
--R
--R Description: Interpret leading open paren as lisp
--R
--R The dalymode option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 16
--S 17 of 75
)set debug dalymode on
--E 17
--S 18 of 75
)set debug dalymode
--R--------------------------- The dalymode Option ---------------------------
--R
--R Description: Interpret leading open paren as lisp
--R
--R The dalymode option may be followed by any one of the following:
--R

--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 18
--S 19 of 75
)set debug
--R                    Current Values of  debug  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rlambdatype   show type information for #1 syntax        on
--Rdalymode     Interpret leading open paren as lisp       on
--R
--E 19
--S 20 of 75
)lisp |$frameAlist|
--R
--RValue = NIL
--E 20
--S 21 of 75
)lisp |$frameNumber|
--R
--RValue = 0
--E 21
--S 22 of 75
)lisp |$currentFrameNum|
--R
--RValue = 0
--E 22
--S 23 of 75
)lisp |$EndServerSession|
--R
--RValue = NIL
--E 23
--S 24 of 75
)lisp |$NeedToSignalSessionManager|
--R
--RValue = T
--E 24
--S 25 of 75
)lisp |$sockBufferLength|

--R
--RValue = 9217
--E 25
--S 26 of 75
)lisp SessionManager
--R
--RValue = 1
--E 26
--S 27 of 75
)lisp |$SessionManager|
--R
--RValue = 1
--E 27
--S 28 of 75
)lisp ViewportServer
--R
--RValue = 2
--E 28
--S 29 of 75
)lisp |$ViewportServer|
--R
--RValue = 2
--E 29
--S 30 of 75
)lisp MenuServer
--R
--RValue = 3
--E 30
--S 31 of 75
)lisp |$MenuServer|
--R
--RValue = 3
--E 31
--S 32 of 75
)lisp SessionIO
--R
--RValue = 4
--E 32
--S 33 of 75
)lisp |$SessionIO|
--R
--RValue = 4

--E 33
--S 34 of 75
)lisp MessageServer
--R
--RValue = 5
--E 34
--S 35 of 75
)lisp |$MessageServer|
--R
--RValue = 5
--E 35
--S 36 of 75
)lisp InterpWindow
--R
--RValue = 6
--E 36
--S 37 of 75
)lisp |$InterpWindow|
--R
--RValue = 6
--E 37
--S 38 of 75
)lisp KillSpad
--R
--RValue = 7
--E 38
--S 39 of 75
)lisp |$KillSpad|
--R
--RValue = 7
--E 39
--S 40 of 75
)lisp DebugWindow
--R
--RValue = 8
--E 40
--S 41 of 75
)lisp |$DebugWindow|
--R
--RValue = 8
--E 41

--S 42 of 75
)lisp NAGLinkServer
--R
--RValue = 8
--E 42
--S 43 of 75
)lisp |$NAGLinkServer|
--R
--RValue = 8
--E 43
--S 44 of 75
)lisp Forker
--R
--RValue = 9
--E 44
--S 45 of 75
)lisp |$Forker|
--R
--RValue = 9
--E 45
--S 46 of 75
)lisp CreateFrame
--R
--RValue = 1
--E 46
--S 47 of 75
)lisp |$CreateFrame|
--R
--RValue = 1
--E 47
--S 48 of 75
)lisp SwitchFrames
--R
--RValue = 2
--E 48
--S 49 of 75
)lisp |$SwitchFrames|
--R
--RValue = 2
--E 49
--S 50 of 75
)lisp EndOfOutput

--S 51 of 75
)lisp |$EndOfOutput|
--R
--RValue = 3
--E 51
--S 52 of 75
)lisp CallInterp
--R
--RValue = 4
--E 52
--S 53 of 75
)lisp |$CallInterp|
--R
--RValue = 4
--E 53
--S 54 of 75
)lisp EndSession
--R
--RValue = 5
--E 54
--S 55 of 75
)lisp |$EndSession|
--R
--RValue = 5
--E 55
--S 56 of 75
)lisp LispCommand
--R
--RValue = 6
--E 56
--S 57 of 75
)lisp |$LispCommand|
--R
--RValue = 6
--E 57
--S 58 of 75
)lisp SpadCommand
--R
--RValue = 7

--E 58
--S 59 of 75
)lisp |$SpadCommand|
--R
--RValue = 7
--E 59
--S 60 of 75
)lisp SendXEventToHyperTeX
--R
--RValue = 8
--E 60
--S 61 of 75
)lisp |$SendXEventToHyperTeX|
--R
--RValue = 8
--E 61
--S 62 of 75
)lisp QuietSpadCommand
--R
--RValue = 9
--E 62
--S 63 of 75
)lisp |$QuietSpadCommand|
--R
--RValue = 9
--E 63
--S 64 of 75
)lisp CloseClient
--R
--RValue = 10
--E 64
--S 65 of 75
)lisp |$CloseClient|
--R
--RValue = 10
--E 65
--S 66 of 75
)lisp QueryClients
--R
--RValue = 11
--E 66

--S 67 of 75
)lisp |$QueryClients|
--R
--RValue = 11
--E 67
--S 68 of 75
)lisp QuerySpad
--R
--RValue = 12
--E 68
--S 69 of 75
)lisp |$QuerySpad|
--R
--RValue = 12
--E 69
--S 70 of 75
)lisp NonSmanSession
--R
--RValue = 13
--E 70
--S 71 of 75
)lisp |$NonSmanSession|
--R
--RValue = 13
--E 71
--S 72 of 75
)lisp KillLispSystem
--R
--RValue = 14
--E 72
--S 73 of 75
)lisp |$KillLispSystem|
--R
--RValue = 14
--E 73
--S 74 of 75
)lisp CreateFrameAnswer
--R
--RValue = 50
--E 74
--S 75 of 75
)lisp |$CreateFrameAnswer|

)spool
)lisp (bye)

———-

References^{[1] nothing}

14